# Opinion Dynamics with Multiple Adversaries

### v1

@@ Abstract.
Opinion dynamics model how the publicly expressed opinions of users in a social network coevolve according to their neighbors as well as their own intrinsic opinion. Motivated by the real-world manipulation of social networks during the 2016 US elections and the 2019 Hong Kong protests, a growing body of work models the effects of a strategic actor who interferes with the network to induce disagreement or polarization. We lift the assumption of a single strategic actor by introducing a model in which any subset of network users can manipulate network outcomes. They do so by acting according to a fictitious intrinsic opinion. Strategic actors can have conflicting goals, and push competing narratives. We characterize the Nash Equilibrium of the resulting meta-game played by the strategic actors. Experiments on real-world social network datasets from Twitter, Reddit, and Political Blogs show that strategic agents can significantly increase polarization and disagreement, as well as increase the “cost” of the equilibrium. To this end, we give worst-case upper bounds on the Price of Misreporting (analogous to the Price of Anarchy). Finally, we give efficient learning algorithms for the platform to (i) detect whether strategic manipulation has occurred, and (ii) learn who the strategic actors are. Our algorithms are accurate on the same real-world datasets, suggesting how platforms can take steps to mitigate the effects of strategic behavior.
conference: ; ;
@@ 1. Introduction
Over the past decade, social media has experienced rapid growth in both usage and significance. Online social networks, which allow users to share updates about their lives and opinions with a broad audience instantaneously, are now utilized by billions of people globally. These platforms serve various purposes, such as being informed about politics, news, health-related updates, products, and many more (Backstrom et al., 2012; Young, 2006; Banerjee et al., 2013; Shearer and Mitchell, 2021).
Unfortunately, networks can induce polarization, with the network connections serving as a pathway for social discord to increase (Musco et al., 2018a; Chen and Rácz, 2021a; Wang and Kleinberg, 2024; Gaitonde et al., 2020). This is a well-studied sociological phenomenon called the filter-bubble theory (Pariser, 2011). The filter-bubble theory argues that personalized algorithms used by online platforms, such as search engines and social media, selectively display content that aligns with a user’s past behaviors, preferences, and beliefs. This customization creates an “invisible algorithmic editing” of the web, isolating individuals within their own ideological bubbles where they encounter only information that reinforces their existing views. As a result, users are less likely to be exposed to diverse perspectives, potentially narrowing their worldview and fostering polarization. Pariser (2011) warns that such bubbles undermine democratic discourse by limiting opportunities for individuals to engage with challenging or unfamiliar ideas.
Additionally, social networks can be manipulated by malicious entities in order to create discord and cause disagreement. For instance, the 2017 indictment of the Russian Internet Research Agency (IRA) by the U.S. Department of Justice Special Counsel’s Office alleged that the IRA leveraged multiple social media accounts and targeted advertising to achieve “a strategic goal to sow discord in the U.S. political system, including the 2016 U.S. presidential election” (Mueller, 2018). In 2019, Twitter, Inc. (2019) disclosed that at least 936 accounts attempted to induce discord in Hong Kong, to e.g. hinder protesters’ ability to organize effectively during the independence movement. As social media continues to proliferate, it is likely that these types of external interferences will become increasingly common. Additionally, networks of Facebook pages have targeted Americans with sports betting scams, amplifying their reach by disseminating provocative conspiracy theories about political figures and natural disasters (Bjork-James and Donovan, 2024). These schemes leverage the economics of the internet, where engagement with inflammatory content is monetized, and social media algorithms inadvertently amplify such content, enabling bad actors to exploit audiences for profit.
To model the opinions’ evolution, computer scientists, sociologists, and statisticians have relied on the framework of opinion dynamics where the users’ opinions coevolve according to a weighted network $G=(V,E,w)$, and each user updates their opinion as a combination of their own intrinsic opinion as well as the opinions of their neighbors (Friedkin and Johnsen, 1990). This model of opinion exchange has the advantage of taking into account both network interactions and their own intrinsic opinion. So far, all of the existing works consider a single actor who has the ability to act on the network to induce disagreement or polarization (Musco et al., 2018a; Chen and Rácz, 2021a; Wang and Kleinberg, 2024; Ristache et al., 2024; Gaitonde et al., 2020; Rácz and Rigobon, 2023; Chitra and Musco, 2020).
@@
For instance, assume a social network where a set of $S$ of political actors want the network to believe that their stance on a topic (e.g., abortion, elections, drug legalization, etc.) is the best. They achieve this by adversarially reporting different intrinsic opinions. This ensures that their influence is both persuasive and credible within the local network context. Such adversarial behavior can result in significantly different (cf. Figure 1) and highly polarized equilibria, where the strategic agents’ opinions appear dominant despite not reflecting the actual intrinsic views of the majority.
@@ 1.1. Our Contributions
Concretely, the true opinions of the agents are $\bm{s}_{1},...,\bm{s}_{n}\in\mathbb{R}$, and there is a set $S$ of deviating agents who report $\{\bm{s}_{i}^{\prime}\}_{i\in S}$. The goal of the strategic agents is to minimize the cost function of Equation 2 at consensus $z^{\prime}=((I-A)L+A)^{-1}A\bm{s}^{\prime}$ where $\bm{s}^{\prime}$ is the vector which has entries $\bm{s}_{i}$ for all $i\notin S$ and $\bm{s}_{i}^{\prime}$ for all $i\in S$. The local optimization of agent $i$ becomes:
@@
We give the Nash equilibrium of the game defined by Equation 3, and show that all Nash-optimal strategies are pure. The Pure Strategy Nash Equilibrium (PSNE) that is given by solving a constrained linear system. Given the PSNE of the game, we characterize the actors who can have the most influence in strategically manipulating the network.
@@ 1.2. Preliminaries and Notations
The set $[n]$ denotes $\{1,\dots,n\}$. $\|\bm{x}\|_{p}$ denotes the $\ell_{p}$-norm of vector $\bm{x}$, whose $i$-th entry is denoted by $\bm{x}_{i}$. $X\succeq 0$ denotes that the matrix $X$ is positive semi-definite and $\|X\|_{2}$ corresponds to the spectral norm of $X$. The Laplacian of the graph $G$ is denoted by $L=D-W$ where $W$ is the weight matrix of the graph, which has entries $w_{ij}\geq 0$, and $D$ is the diagonal degree matrix with diagonal entries $D_{ii}=\sum_{i\sim j}w_{ij}$. The Laplacian has eigenvalues $0=\lambda_{1}\leq\lambda_{2}\leq\dots\leq\lambda_{n}$. For any undirected and connected graph $G$, $L$ is symmetric and PSD, so we can write the eigendecomposition of $L$ as:
@@
(6) Polarization $\textrm{Polarization Ratio}\qquad\mathcal{P}(\bm{z})$ $=\sum_{i\in[n]}(\bm{z}_{i}-\bar{z})^{2},\text{ where }\bar{z}=\frac{1}{n}\sum_{i\in[n]}\bm{z}_{i},$
(7) Disagreement $\textrm{Disagreement Ratio}\qquad\mathcal{D}(\bm{z})$ $=\sum_{i,j\in[n]}w_{ij}(\bm{z}_{i}-\bm{z}_{j})^{2}=\bm{z}^{T}L\bm{z}.$
@@
Unlike the Price of Anarchy (PoA), the equilibrium $\bm{z}$ in the denominator of Eq. (8) is the Nash equilibrium for the Friedkin-Johnson dynamics without manipulation. In the PoA, the denominator would be $C(\bm{z}^{*})$, where $\bm{z}^{*}$ is a socially optimal equilibrium (Bindel et al., 2011). Since we study strategic manipulations as a meta-game with respect to the base game of FJ dynamics, it is more relevant for us to compare $\bm{z}^{\prime}$ with $\bm{z}$ than with $\bm{z}^{*}$. Note that $C(\bm{z}^{*})\leq C(\bm{z})\leq C(\bm{z}^{\prime})$, so $\mathsf{PoA}\geq\mathsf{PoM}\geq 1$ always.
@@ 1.3. Related Work
[…]
@@
Motivated by real-world manipulation of social networks in, e.g., the 2016 US election, a recent line of work studies polarization and strategic behavior in opinion dynamics (Gaitonde et al., 2020, 2021; Chen and Rácz, 2021b; Wang and Kleinberg, 2024; Ristache et al., 2024, 2025). Chen and Rácz (2021b) consider a model in which an adversary can control $k\leq n$ nodes’ internal opinions and seeks to maximize polarization at equilibrium. Similarly, Gaitonde et al. (2020) considers a single adversary who can modify intrinsic opinions $\bm{s}$ belonging to an $\ell_{2}$-ball. More recent work also studies modification of agents’ susceptibility parameters $\alpha_{i}$ to alter the median opinion at equilibrium (Ristache et al., 2025). By contrast, we study a setting in which any subset $S\subseteq[n]$ can be strategic. Unlike previous works, these “adversaries” can have conflicting goals in our model.
@@
1.4. Real-world Datasets
To support our results, we use data grounded in practice, which have also been used in previous studies to study polarization and disagreement (cf. Musco et al. (2018b); Chitra and Musco (2020); Wang and Kleinberg (2024); Adamic and Glance (2005)). Specifically, we use Twitter, Reddit, and Political blog networks, summarized in Table 1 summarizes these. Both the Twitter and Reddit datasets are due to Chitra and Musco (2020). The vectors $\bm{s}$ of initial opinions for both are obtained via sentiment analysis and also follow the post-processing of Wang and Kleinberg (2024).
(1) Twitter dataset.
@@ 1.4. Real-world Datasets
(3) Political Blogs (Polblogs) dataset.
@@
Network Nodes ($n$) Edges ($m$) Description
Twitter 548 3,638 User interactions during 2013 Delhi elections.
Reddit 556 8,969 User interactions in r/politics subreddit
Polblogs 1,490 16,178 Liberal and conservative blog network
Table 1. Summary of the social network datasets we use.
@@ 2. Strategic Opinion Formation
The opinion formation game has two phases. First, strategic agents privately choose a strategic intrinsic opinion according to Equation 3. Second, agents exchange opinions and reach consensus as if they were in the Friedkin-Johnson dynamics, except the strategic opinions are used in place of the true intrinsic opinions.
@@
Let $S\subseteq[n]$ be a set of strategic agents. If agent $i\in S$ wants network members to express opinions close to $\bm{s}_{i}$, what choice of $\bm{s}_{i}^{\prime}$ is optimal and minimizes the cost function of Equation 3?
The following theorem characterizes the Nash Equilibria of the Intrinsic Belief Lying Problem.
@@
Proof.
$\nabla_{\bm{s}^{\prime}}c_{i}(\bm{s}^{\prime})$ $=(1-\alpha_{i})\sum_{j\sim i}w_{ij}2(B^{T}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{T}B)(\bm{s}^{\prime})+\alpha_{i}(2B^{T}\bm{e}_{i}\bm{e}_{i}^{T}B\bm{s}^{\prime}-2B^{T}\bm{e}_{i}\bm{e}_{i}^{T}\bm{s}),$
$\nabla_{\bm{s}^{\prime}}^{2}c_{i}(\bm{s}^{\prime})$ $=2(1-\alpha_{i})B^{T}\bigg{[}\sum_{j\sim i}w_{ij}2(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{T}\bigg{]}B+2\alpha_{i}B^{T}\bm{e}_{i}\bm{e}_{i}^{T}B.$
Let $L_{i}\in\mathbb{R}^{n\times n}$ be:
$L_{i}:=\sum_{j\sim i}w_{ij}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{T}.$
00 0 $=\frac{1}{2}\frac{\partial}{\partial\bm{s}_{i}^{\prime}}c_{i}(\bm{s}^{\prime})$
$=\bm{e}_{i}^{T}(1-\alpha_{i})B^{T}\bigg{[}\sum_{j\sim i}w_{ij}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{T}\bigg{]}B\bm{s}^{\prime}+\bm{e}_{i}^{T}\alpha_{i}(B^{T}\bm{e}_{i}\bm{e}_{i}^{T}Bs^{\prime}-B^{T}\bm{e}_{i}\bm{e}_{i}^{T}\bm{s})$
The above display gives the solution for $\bm{s}_{i}^{\prime}$ in terms of all entries of $s^{\prime}$. Assembling the critical points into a linear system, we obtain precisely that for all $i\in S$, $\bm{e}_{i}^{T}\mathbb{T}_{i}\bm{s}^{\prime}=\bm{y}_{i}$. Since $\bm{s}_{j}^{\prime}=\bm{s}_{j}$ for $j\not\in S$, the overall linear system describes the Nash equilibria. ∎
To illustrate the Theorem, we consider a toy example.
Example 2.3 (Two-Node Graph).
Consider a graph with 2 nodes and one edge with weight $w>0$. We set $\alpha_{1}=\alpha_{2}=0.5$ for simplicity. Suppose that both agents deviate, i.e., $S=[2]$. Then, we can calculate $B$ to be
$B=\frac{1}{2w+1}\begin{pmatrix}w+1&w\\ w&w+1\end{pmatrix}$
and
(9) $\bm{z}_{0}^{\prime}=\frac{(w+1)\bm{s}_{0}^{\prime}+w\bm{s}_{1}^{\prime}}{2w+1},\quad\bm{z}_{1}^{\prime}=\frac{w\bm{s}_{0}^{\prime}+(w+1)\bm{s}_{1}^{\prime}}{2w+1},$
yielding the two cost functions
$c_{0}(\bm{s}^{\prime})$ $=\frac{1}{2}w\left(\frac{\bm{s}_{0}^{\prime}-\bm{s}_{1}^{\prime}}{2w+1}\right)^{2}+\frac{1}{2}\left(\frac{(w+1)\bm{s}_{0}^{\prime}+w\bm{s}_{1}^{\prime}}{2w+1}-\bm{s}_{0}\right)^{2}$
$c_{1}(\bm{s}^{\prime})$ $=\frac{1}{2}w\left(\frac{\bm{s}_{0}^{\prime}-\bm{s}_{1}^{\prime}}{2w+1}\right)^{2}+\frac{1}{2}\left(\frac{(w+1)\bm{s}_{1}^{\prime}+w\bm{s}_{0}^{\prime}}{2w+1}-\bm{s}_{1}\right)^{2}.$
Taking the first order conditions $\frac{\partial c_{0}}{\partial\bm{s}_{0}^{\prime}}=0$ and $\frac{\partial c_{1}}{\partial\bm{s}_{1}^{\prime}}=0$ we get a linear system whose solutions are:
$\bm{s}_{0}^{\prime}$ $=\frac{w^{2}(s_{0}-s_{1})+(3w+1)\bm{s}_{0}}{3w+1},\quad\bm{s}_{1}^{\prime}=\frac{w^{2}(\bm{s}_{0}-\bm{s}_{1})+(3w+1)s_{1}}{3w+1}.$
Replacing these values back to the costs we get that
$\forall i:c_{i}(s_{0}^{\prime},s_{1}^{\prime})=\frac{1}{2}\frac{w(w^{2}+3w-1)(\bm{s}_{0}-\bm{s}_{1})^{2}}{9w^{2}+6w+1},$
On the other hand, if all agents are honest, then the cost for each is:
$\forall i:c_{i}(\bm{s}_{0},\bm{s}_{1})=\frac{1}{2}\frac{w(w+1)(\bm{s}_{0}-\bm{s}_{1})^{2}}{(2w+1)^{2}}.$
and the ratio of the two costs is at least $\max\{1,w/3\}$.
@@
Let $T\in\mathbb{R}^{|S|\times n}$ have rows $\{\bm{e}_{i}^{T}\mathbb{T}_{i}\}_{i\in S}$ given by Theorem 2.2. Let $\tilde{T}\in\mathbb{R}^{|S|\times|S|}$ be the submatrix of $T$ selecting columns belonging to $S$. Let $\bm{y}\in\mathbb{R}^{|S|}$ have entries $\bm{y}_{i}=\alpha_{i}B_{ii}\bm{s}_{i}$ as above. Let $\tilde{\bm{y}}=\bm{y}-\sum_{j\not\in S}\bm{s}_{j}T\bm{e}_{j}$. Then the set of Nash equilibria, if any exist, are given by the solutions to the unconstrained linear system
@@
Thus, in a Nash equilibrium, every strategic agent solves their corresponding equation given by Equation 10. The explicit characterization of equilibria also implies that Nash equilibria cannot be mixed.
@@
The Nash equilibrium corresponds to solving the system of $|S|$ linear equations in the scalars $\{\bm{s}_{i}^{\prime}|i\in S\}$ given by Equation 10. Also, all Nash equilibria are pure-strategy Nash equilibria.

### v2

@@ Abstract.
Opinion dynamics models how the publicly expressed opinions of users in a social network coevolve according to their neighbors as well as their own intrinsic opinion. Motivated by the real-world manipulation of social networks during the 2016 US elections and the 2019 Hong Kong protests, a growing body of work models the effects of a strategic actor who interferes with the network to induce disagreement or polarization. We lift the assumption of a single strategic actor by introducing a model in which any subset of network users can manipulate network outcomes. They do so by acting according to a fictitious intrinsic opinion. Strategic actors can have conflicting goals, and push competing narratives. We characterize the Nash Equilibrium of the resulting meta-game played by the strategic actors. Experiments on real-world social network datasets from Twitter, Reddit, and Political Blogs show that strategic agents can significantly increase polarization and disagreement, as well as increase the “cost” of the equilibrium. To this end, we give worst-case upper bounds on the Price of Misreporting (analogous to the Price of Anarchy). Finally, we give efficient learning algorithms for the platform to (i) detect whether strategic manipulation has occurred, and (ii) learn who the strategic actors are. Our algorithms are accurate on the same real-world datasets, suggesting how platforms can take steps to mitigate the effects of strategic behavior.
platform design; misreporting; social networks
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3792080isbn: 979-8-4007-2307-0/2026/04ccs: Information systems World Wide Webccs: Information systems Social networksccs: Information systems Electronic commerce
@@ 1. Introduction
Over the past decade, social media platforms have grown dramatically in both scale and societal influence, enabling billions of users to share information and opinions instantaneously. Online social networks now play a central role in how individuals consume news, form political views, and engage with issues related to health, products, and public life (Backstrom et al., 2012; Young, 2006; Banerjee et al., 2013; Shearer and Mitchell, 2021).
A growing body of evidence suggests, however, that these platforms can exacerbate polarization by shaping and amplifying patterns of social interaction (Musco et al., 2018a; Chen and Rácz, 2021a; Wang and Kleinberg, 2024; Gaitonde et al., 2020). One prominent explanation is the filter-bubble theory (Pariser, 2011), whereby personalized algorithms selectively expose users to content aligned with their past behavior and beliefs. This form of “invisible algorithmic editing” limits exposure to diverse viewpoints, reinforcing existing opinions and potentially undermining democratic discourse by reducing engagement with challenging or unfamiliar ideas.
Beyond endogenous polarization, social networks are increasingly subject to strategic manipulation by malicious actors seeking to induce discord. Notable examples include the alleged interference of the Russian Internet Research Agency in the 2016 U.S. presidential election (Mueller, 2018), coordinated attempts to disrupt protest organizations in Hong Kong (Twitter, Inc., 2019), and large-scale scams exploiting political and conspiratorial content to maximize engagement and profit (Bjork-James and Donovan, 2024). These interventions leverage the economics of attention on digital platforms, where inflammatory content is rewarded and algorithmically amplified, making social networks particularly vulnerable to strategic misuse.
To study the evolution of beliefs in such settings, researchers across computer science, sociology, and statistics have relied on models of opinion dynamics, in which agents’ expressed opinions coevolve over a weighted network by combining intrinsic beliefs with social influence from neighbors (Friedkin and Johnsen, 1990). While this framework captures the interplay between individual predispositions and network effects, existing work largely focuses on scenarios involving a single adversarial or coordinating actor acting on the network to induce polarization or disagreement (Musco et al., 2018a; Chen and Rácz, 2021a; Wang and Kleinberg, 2024; Ristache et al., 2024; Gaitonde et al., 2020; Rácz and Rigobon, 2023; Chitra and Musco, 2020).
@@
For instance, assume a social network where a set of $S$ of political actors want the network to believe that their stance on a topic (e.g., abortion, elections, drug legalization, etc.) is the best. They achieve this by adversarially reporting different intrinsic opinions. This ensures that their influence is both persuasive and credible within the local network context. Such adversarial behavior can result in significantly different (cf. Figure˜1) and highly polarized equilibria, where the strategic agents’ opinions appear dominant despite not reflecting the actual intrinsic views of the majority.
@@
Concretely, the true opinions of the agents are $\bm{s}_{1},...,\bm{s}_{n}\in\mathbb{R}$, and there is a set $S$ of deviating agents who report $\{\bm{s}_{i}^{\prime}\}_{i\in S}$. While an agent $i\in S$ has a true intrinsic opinion $\bm{s}_{i}$, she may report some $\bm{s}_{i}^{\prime}\neq\bm{s}_{i}$ for the purpose of manipulating the dynamics of Equation˜1. Specifically, her goal is to minimize her unique cost function, given by Equation˜2, at the resulting consensus $z^{\prime}=((I-A)L+A)^{-1}A\bm{s}^{\prime}$ where $\bm{s}^{\prime}$ is the vector which has entries $\bm{s}_{i}$ for all $i\notin S$ and $\bm{s}_{i}^{\prime}$ for all $i\in S$. The local optimization of agent $i$ becomes:
@@
Note that while Equation˜3 depends on the choices of all $j\in S$, agent $i$ can only choose the $i^{th}$ entry of $\bm{s}^{\prime}$.
@@
We give the Nash equilibrium of the game defined by Equation˜3, and show that all Nash-optimal strategies are pure. The Pure Strategy Nash Equilibrium (PSNE) that is given by solving a constrained linear system. Given the PSNE of the game, we characterize the actors who can have the most influence in strategically manipulating the network.
@@ 1.1. Preliminaries and Notations
The set $[n]$ denotes $\{1,\dots,n\}$. $\|\bm{x}\|_{p}$ denotes the $\ell_{p}$-norm of vector $\bm{x}$, whose $i$-th entry is denoted by $\bm{x}_{i}$. $X\succeq 0$ denotes that the matrix $X$ is positive semi-definite and $\|X\|_{2}$ corresponds to the spectral norm of $X$. The Laplacian of the graph $G$ is denoted by L=DWL=D-W where $W$ is the weight matrix of the graph, which has entries $w_{ij}\geq 0$, and $D$ is the diagonal degree matrix with diagonal entries $D_{ii}=\sum_{i\sim j}w_{ij}$. The Laplacian has eigenvalues $0=\lambda_{1}\leq\lambda_{2}\leq\dots\leq\lambda_{n}$. For any undirected and connected graph $G$, $L$ is symmetric and PSD, so we can write the eigendecomposition of $L$ as:
@@
(6) $\textrm{Polarization}\qquad\mathcal{P}(\bm{z})$ $=\sum_{i\in[n]}(\bm{z}_{i}-\bar{z})^{2},\bar{z}=\frac{1}{n}\sum_{i\in[n]}\bm{z}_{i},$
(7) $\textrm{Disagreement}\qquad\mathcal{D}(\bm{z})$ $=\sum_{i,j\in[n]}w_{ij}(\bm{z}_{i}-\bm{z}_{j})^{2}=\bm{z}^{T}L\bm{z}.$
@@
Unlike the Price of Anarchy (PoA), the equilibrium $\bm{z}$ in the denominator of Eq. (8) is the Nash equilibrium for the Friedkin-Johnson dynamics without manipulation. In the PoA, the denominator would be $C(\bm{z}^{*})$, where $\bm{z}^{*}$ is a socially optimal equilibrium (Bindel et al., 2011). Since we study strategic manipulations as a meta-game with respect to the base game of FJ dynamics, it is more relevant for us to compare $\bm{z}^{\prime}$ with $\bm{z}$ than with $\bm{z}^{*}$.
@@ 1.2. Related Work
Opinion dynamics are well-studied in computer science and economics, as well as sociology, political science, and related fields. There have been many models proposed for opinion dynamics, such as with network interactions as we study in this paper (FJ model) (Friedkin and Johnsen, 1990; Bindel et al., 2015), bounded confidence dynamics (Hegselman-Krausse Model) (Hegselmann et al., 2002), coevolutionary dynamics (Bhawalkar et al., 2013) as well as many variants of them; see, for example Abebe et al. (2018); Hązła et al. (2019); Fotakis et al. (2016, 2023); Ristache et al. (2024). The work of (Bindel et al., 2011) shows bounds on the Price of Anarchy (PoA) between the PSNE and the welfare-optimal solution for the FJ model, and the subsequent work of (Bhawalkar et al., 2013) shows PoA bounds for the coevolutionary dynamics. Additionally, the opinion dynamics have been modeled by the control community; see, for example, (Nedić and Touri, 2012; De Pasquale and Valcher, 2022; Bhattacharyya et al., 2013; Chazelle, 2011).
@@
Motivated by real-world manipulation of social networks in, e.g., the 2016 US election, a recent line of work studies polarization and strategic behavior in opinion dynamics (Gaitonde et al., 2020, 2021; Chen and Rácz, 2021b; Wang and Kleinberg, 2024; Ristache et al., 2024, 2025). Chen and Rácz (2021b) consider a model in which an adversary can control $k\leq n$ nodes’ internal opinions and seeks to maximize polarization at equilibrium. Similarly, Gaitonde et al. (2020) considers a single adversary who can modify intrinsic opinions $\bm{s}$ belonging to an $\ell_{2}$-ball. More recent work also studies modification of agents’ susceptibility parameters $\alpha_{i}$ to alter the median opinion at equilibrium (Ristache et al., 2025). Our work shares a similar perspective to these: non-strategic agents follow the base-game (in our case, Friedkin-Johnson dynamics Equation˜1), while strategic agents treat one or more variables in the base game as a strategic choice that they may vary to manipulate game outcomes (in our case, the misreported intrinsic opinions $\{\bm{s}_{i}^{\prime}\}_{i\in S}$). This perspective emphasizes the dynamics and equilibria of the resulting meta-game, which is played by the strategic agents against the backdrop of the base-game dynamics. Unlike these previous works, however, in our setting any subset $S\subseteq[n]$ can be strategic, and these “adversaries” can have conflicting goals.
@@
Real-world Datasets
To support our results, we use data grounded in practice, which have also been used in previous studies to study polarization and disagreement (cf. Musco et al. (2018b); Chitra and Musco (2020); Wang and Kleinberg (2024); Adamic and Glance (2005)). Specifically, we use Twitter, Reddit, and Political blog networks. Both the Twitter and Reddit datasets are due to Chitra and Musco (2020). The vectors $\bm{s}$ of initial opinions for both are obtained via sentiment analysis and also follow the post-processing of Wang and Kleinberg (2024).
(1) Twitter dataset ($n=548,m=3638$).
@@
(3) Political Blogs (Polblogs) dataset ($n=1490,m=16178$).
@@ 2. Strategic Opinion Formation
The opinion formation game has two phases. First, strategic agents privately choose a strategic intrinsic opinion according to Equation˜3. Second, agents exchange opinions and reach consensus as if they were in the Friedkin-Johnson dynamics, except the strategic opinions are used in place of the true intrinsic opinions.
@@
Let $S\subseteq[n]$ be a set of strategic agents. If agent $i\in S$ wants network members to express opinions close to $\bm{s}_{i}$, what choice of $\bm{s}_{i}^{\prime}$ is optimal and minimizes the cost function of Equation˜3?
The following theorem characterizes the Nash Equilibria of the Intrinsic Belief Lying Problem (cf. Appendix˜A for a proof).
@@
Let $T\in\mathbb{R}^{|S|\times n}$ have rows $\{\bm{e}_{i}^{T}\mathbb{T}_{i}\}_{i\in S}$ given by Theorem˜2.2. Let $\tilde{T}\in\mathbb{R}^{|S|\times|S|}$ be the submatrix of $T$ selecting columns belonging to $S$. Let $\bm{y}\in\mathbb{R}^{|S|}$ have entries $\bm{y}_{i}=\alpha_{i}B_{ii}\bm{s}_{i}$ as above. Let $\tilde{\bm{y}}=\bm{y}-\sum_{j\not\in S}\bm{s}_{j}T\bm{e}_{j}$. Then the set of Nash equilibria, if any exist, are given by the solutions to the unconstrained linear system
@@
Thus, in a Nash equilibrium, every strategic agent solves their corresponding equation given by Equation˜9. The explicit characterization of equilibria also implies that Nash equilibria cannot be mixed.
@@
The Nash equilibrium corresponds to solving the system of $|S|$ linear equations in the scalars $\{\bm{s}_{i}^{\prime}|i\in S\}$ given by Equation˜9. Also, all Nash equilibria are pure-strategy Nash equilibria.